two resistance 60.36 and 30.09 are connected in parllel. the equivalent resistance is 

Grade:12th pass

1 Answers

Muhammmad Hamza Faryad
2 years ago
Here R1= 60.36 and R2 = 30.09, Both are connected in parallel.
Their equivalent resistance will be ‘Req
Req = R1.R2/R1+R= (60.36)(30.09) / 60.36 + 30.09
Req = 1816.2324 / 90.45 =  20. 079 ohm